Inverse Scattering Theory for Jacobi Operators with Quasi-Periodic Background

Gerald Teschl University of Vienna, Austria

We will consider inverse scattering theory for Jacobi operators which are short range perturbation of a quasi-periodic background operators. Neccessary and sufficcient conditions for given scattering data to determine a unique Jacobi operator are presented.